A train travels north at 100 kilometers per hour and travels for 160mins. Then it turns east and travels 2000 meters per minute for 2 hours. How far has the train traveled?
Answer by stanbon(75887)

A train travels north at 100 kilometers per hour and travels for 160mins. Then it turns east and travels 2000 meters per minute for 2 hours. How far has the train traveled?
-------------------------
North segment DATA:
rate = 100 km/hr ; time = (160/60) = 8/3 hrs ; distance = (800/3)km
--------------------------------------
East segment DATA:
rate = 2000 meter/minute ; time = 120 minutes ; distance = 240,000 km
----------------------------------
Total distance traveled:
(800/3) km + (240,000) km = 240.266.67 km
===========================================
Cheers,
Stan H.
=================